# Quadrotor Navigation using Reinforcement Learning with Privileged Information

### v1

@@ I INTRODUCTION
To bridge these gaps in the state of the art, we extend the method of Zhang et al. [2] to enable navigation around large obstacles while retaining the efficient differentiable simulation-based training pipeline. We provide the following contributions:
@@
    an approach to reduce attitude control response latency by computing angular rate setpoints from predicted actions;
@@
    an approach to mitigate the effects of inaccurate motor and inertial parameter estimation by leveraging domain randomization; and
  5. 5.
    extensive evaluation of the system in photo-realistic simulation and hardware experiments [^1] A video of the experiments may be found at https://youtu.be/MHDgV55PFOo. as well as an open-source software release [^2] https://github.com/rislab/depthnav.
@@ II RELATED WORK
In classical global path planning, occupancy grids or Euclidean signed distance fields (ESDF) are used to avoid getting stuck in local extrema. In contrast, few learning-based navigation methods exploit such maps as privileged information available during training but absent at inference [6, 3]. YOPO [3] trains a neural policy to generate motion primitive offsets and scores using an ESDF cost map as supervision. However, this approach provides only limited guidance for escaping large convex obstacle regions.
Our approach improves upon the work of Zhang et al. [2] by incorporating a yaw prediction objective to navigate in complex environments that require sudden changes in orientation (e.g, sharp corners, walls, dead ends, etc.). In contrast to [3], we enable robustness to local extrema via privileged information about ToA maps to goal positions without using a high-cost training simulator. In addition, both [2, 3] assume perfect system knowledge for zero-shot sim-to-real transfer. Conversely, our approach ensures robustness to incorrect motor or inertial parameter identification via domain randomization during training.
@@ III-B Network Architecture
The state consists of the current velocity and orientation of the robot. The target is defined by a desired velocity vector towards the goal and a goal importance value, defined as the reciprocal of the goal distance, which provides contextual information about the goal’s proximity relative to obstacles in the depth image, helping the network distinguish whether the goal lies in front of or behind an obstacle. All input vectors are expressed with respect to the starting frame of the robot. Our network differs from that of [2] in that we use a separate target extractor which includes a goal importance value, we use normalization layers to balance features, and we predict a yaw angle in addition to mass-normalized thrust.
@@ III-C3 Target Velocity Loss
$L_{\text{vmax}}$ $=\frac{1}{T}\sum_{k=1}^{T}\max(\lVert\mathbf{v}_{k}\rVert-v_{\max})^{2}$ (10)
@@ III-C5 Privileged Information via ToA Map
We observe that existing obstacle avoidance and collision loss terms in prior work [2] are often insufficient for navigation around large obstacles. Local collision losses penalize proximity or approach speed to surfaces but do not actively encourage progress or escape, which can cause the robot to become stuck in concave obstacle regions. To address these limitations, we use time-of-arrival (ToA) maps, whose gradients guide the robot’s velocity toward the goal while avoiding obstacles (see Fig. 4).
@@
$r$ is the robot radius, $d(\mathbf{x})$ is the distance to the nearest obstacle, $d_{\text{safe}}$ is the threshold distance considered safe, and $v_{\text{slow}}$ is the minimum travel speed near an obstacle. Points at least $d_{\text{safe}}$ from obstacles allow the wavefront to travel at $1\text{\,}\mathrm{m}\text{/}\mathrm{s}$, while points closer to obstacles have reduced speed proportional to $d(\mathbf{x})$, which biases paths away from surfaces (Fig. 4b).
$T(\mathbf{x})$ is computed on a volumetric grid using the Fast Marching Method (FMM) [10] via the scikit-fmm library [^3] https://github.com/scikit-fmm/scikit-fmm. Gradients $\nabla T(\mathbf{x})$ are calculated at grid cell centers via finite differencing and interpolated to provide values in continuous space. The negative gradient then defines the velocity set-point direction $\mathbf{v}_{k}^{\text{set}}$ in the target velocity loss (see Eq. 9), providing a consistent training signal that encourages the robot to make progress toward the goal while avoiding obstacles.
@@ III-D Training
Figure 6: Comparison of attitude control performance without LABEL:sub@sfig:without_omega and with LABEL:sub@sfig:with_omega the derivative feedback term, $\boldsymbol{\omega}_{d}$, in the attitude controller.
@@ III-E Body Rate Attitude Control
To bridge the sim-to-real gap between point-mass dynamics in simulation and rigid-body dynamics in reality, we adopt a PD attitude controller [13] that tracks desired thrust $\mathbf{F}$, orientation $\mathbf{R}_{d}$, and body rates $\boldsymbol{\omega}_{d}$.
The desired rotation matrix $\mathbf{R}_{d}$ is constructed such that the body $\mathbf{z}^{B}$ axis aligns with the predicted thrust vector $\mathbf{F}$ and $\mathbf{x}^{B}$ aligns with the projection of the desired yaw direction $\psi_{d}$ onto the plane orthogonal to $\mathbf{z}^{B}$.
The desired body rates $\boldsymbol{\omega}_{d}$ are obtained by first estimating the Euler angle rates from two consecutive desired attitudes, $\mathbf{R}_{d1}$ and $\mathbf{R}_{d2}$. The relative rotation between them yields the ZYX Euler angle derivatives $\dot{\phi},\dot{\theta},\dot{\psi}$, which are then converted to body rates via the Euler angle Jacobian:
@@
Figure 8: Trajectory profiles overlaid on ground-truth point clouds, with a cross-section shown for clarity. Trajectories are colored by speed (red to yellow, up to $v_{\text{max}}=3\ \mathrm{m/s}$), with body frames every 2 seconds. Start and goal are marked in blue and magenta, respectively. Successful trials are marked with a  in Industry 2 (top) and Cave 2 (bottom) environments.
@@ IV-B Diverse Simulated Environments
We evaluate our planner in both simulated indoor and outdoor environments featuring narrow corridors and large obstacles. The evaluation uses the 11 out-of-distribution scenarios from [14], leveraging a full rigid-body quadrotor dynamics model and depth images rendered with Flightmare [15]. 1,350 simulation trials are conducted across all environments and different start-end location pairs. Further details on the experimental design are available in [14].
We compare our method against two baselines: Back to Newton’s Laws (BNL) [2], trained using the authors’ public codebase444 https://github.com/HenryHuYu/DiffPhysDrone , and an ablated version of our model trained without the ToA privileged information (yaw w/o ToA). All approaches are trained using only primitive obstacles and point-mass dynamics. Performance is measured by success rate and failure modes (collisions and timeouts), with success defined as reaching the target without collision within a fixed time limit.
Our method achieves a 86% success rate, 36% higher than the baselines, and has the lowest collision rate across all approaches (Fig. 8). As shown in Fig. 8, BNL struggles with large, flat obstacles that require turning or lateral movement, often resulting in collisions. The ablated model, lacking geometric awareness, frequently gets stuck in cluttered scenes. In contrast, our full model learns to yaw around large obstacles and leverage ToA map cues to identify collision-free paths through free space.
@@ V HARDWARE EXPERIMENTS
Figure 9: Hardware ablation experiments studying the effect of training a policy without gravity randomization LABEL:sub@fig:no_gravity_compensation and the performance improvement LABEL:sub@fig:gravity_compensation when gravity randomization is applied. The top row shows snapshots of the quadrotor in flight with the red line indicating the goal setpoint. In LABEL:sub@fig:no_gravity_compensation, the start and goal positions coincide, but upon entering autonomous mode, the vehicle quickly loses altitude. In contrast, when domain randomization is used LABEL:sub@fig:gravity_compensation, the robot gains altitude and reaches the goal setpoint. The middle row plots the quadrotor’s z-height measured by a motion capture system. With domain randomization, altitude error is significantly reduced compared to the non-randomized case. The bottom row shows the normalized thrust predicted by the policy, with the domain randomized policy producing an initial thrust of $1.3g$, substantially higher than in the non-randomized case, enabling better compensation for platform modeling inaccuracies.
A custom quadrotor platform was developed to support hardware experimentation. The vehicle spans $15\text{\,}\mathrm{cm}\text{/}$ from rotor to rotor and is equipped with a forward-facing Intel RealSense D456 depth camera, a downward-facing LW20 range finder, and a Matrix Vision BlueFox2 global shutter greyscale camera. All onboard computation is performed by an NVIDIA Orin NX module with 16 GB of RAM. The system has a total mass of $1.7\text{\,}\mathrm{kg}\text{/}$. A TBS Lucid H7 flight controller runs custom Betaflight firmware555 https://github.com/betaflight/betaflight to provide IMU data at $1000\text{\,}\mathrm{Hz}\text{/}$. For GPS-denied state estimation, we employ a monocular visual-inertial navigation system developed by Yao [16]. The RealSense camera generates depth images at $60\text{\,}\mathrm{Hz}\text{/}$ with a resolution of 640 $\times$ 480 pixels and the policy runs at $50\text{\,}\mathrm{Hz}\text{/}$.
@@ V-B Outdoor Navigation Experiments
Figure 10: Outdoor obstacle avoidance test under tree canopy (Table III Forest Flight 4). The policy predicts up to $30\text{\,}\mathrm{\SIUnitSymbolDegree}\text{/}$ in yaw to navigate through dense underbrush with speeds up to $3.8\text{\,}\mathrm{m}\text{/}\mathrm{s}$. From top to bottom: VINS trajectory overlaid on terrain map, onboard RGB images, inverted and max pooled depth images, and velocity profile.
@@
Outdoor Flight Arena
1 14.4 12.1 2.0 1.5 0.7
@@ VI CONCLUSION
We developed a simulation and training framework for efficiently learning vision-based navigation policies that map depth and state observations to thrust and heading commands. Trained using simplified point-mass dynamics and privileged information, the policy generalizes to full quadrotor dynamics and reliably navigates through cluttered environments. Domain randomization, especially over gravity, enables robust feedback control, correcting for modeling errors such as a 15% thrust mismatch. Our approach achieves an 86% success rate in simulation and completes 20 hardware flights covering $589\text{\,}\mathrm{m}\text{/}$ without collisions.
While an improvement over the baselines, the proposed method has difficulty with backtracking in dead ends and exhibits initial yaw oscillations. A promising direction for future work is to explore more expressive learned-memory architectures to enhance spatial reasoning and long-horizon planning without relying on explicit maps. Additionally, incorporating task-specific inputs and new objective functions could extend the method’s use to a wider range of applications.
ACKNOWLEDGEMENTS

### v2

@@ I INTRODUCTION
To bridge these gaps in the state of the art, we extend the method of Zhang et al. [2] to enable navigation around large obstacles. Like Zhang et al. [2], our policy is reactive at deployment, relying only on depth observations and state estimates. However, by training with ToA maps as privileged information, our policy learns globally-aware navigation behavior without requiring a ToA map at test time. We provide the following contributions:
@@
    approaches to bridge the sim-to-real gap via body rate attitude control and domain randomization; and
@@
    extensive evaluation of the system in photo-realistic simulation and hardware experiments [^1] A video of the experiments may be found at https://youtu.be/RbHJ69o-zUc. as well as an open-source software release [^2] https://github.com/rislab/depthnav.
@@ II RELATED WORK
Recent learning-based navigation methods exploit geometric maps as privileged information available during training but absent at inference [6, 3]. YOPO [3] trains a neural policy to generate motion primitive offsets and scores using a Euclidean signed distance field (ESDF) cost map as supervision. However, this approach provides only limited guidance for escaping large convex obstacle regions.
Our approach addresses two limitations of prior work. First, Zhang et al. [2] maintains a fixed heading toward the target, limiting navigation around large obstacles. Our yaw alignment loss directly supervises heading prediction, enabling the robot to yaw while changing direction. Second, as ESDFs encode obstacle distance rather than path direction, Lu et al. [3] provides local avoidance but not globally optimal routing. Our ToA maps address this by guiding the robot along shortest paths. Together, the yaw alignment loss enables reorientation while the ToA gradient teaches the robot which direction to head to make progress towards the goal.
@@ III-B Network Architecture
The state consists of the current velocity and orientation of the robot. The target is defined by a desired velocity vector towards the goal and a goal importance value, defined as the reciprocal of the goal distance. This importance value provides contextual information about the goal’s proximity relative to obstacles in the depth image, helping the network distinguish whether the goal lies in front of or behind an obstacle. All input vectors are expressed with respect to the starting frame of the robot. Our network differs from that of [2] in that we use a separate target extractor which includes a goal importance value, we use normalization layers to balance features, and we predict a yaw angle in addition to mass-normalized thrust.
@@ III-C3 Target Velocity Loss
$L_{\text{vmax}}$ $=\frac{1}{T}\sum_{k=1}^{T}\max(\lVert\mathbf{v}_{k}\rVert-v_{\max},0)^{2}$ (10)
@@ III-C5 Privileged Information via ToA Map
Existing obstacle avoidance and collision loss terms in prior work [2] are insufficient for navigation around large obstacles. Local collision losses penalize proximity or approach speed to surfaces but do not actively encourage progress or escape, which can cause the robot to become stuck in concave obstacle regions. To address these limitations, we use time-of-arrival (ToA) maps, whose gradients guide the robot’s velocity toward the goal while avoiding obstacles (see Fig. 4). Through the ToA loss, the policy learns to infer optimal navigation directions from depth alone, requiring no map at inference.
@@
$r$ is the robot radius, $d(\mathbf{x})$ is the distance to the nearest obstacle, $d_{\text{safe}}$ is the threshold distance considered safe, and $v_{\text{slow}}$ is the minimum travel speed near an obstacle, biasing paths away from surfaces (Fig. 4b).
$T(\mathbf{x})$ is computed on a volumetric grid using the Fast Marching Method (FMM) [10] via the scikit-fmm library [^3] https://github.com/scikit-fmm/scikit-fmm. FMM runs in $\mathcal{O}(N\log N)$ time, where $N$ is the number of grid cells, and each ToA map is pre-computed offline once per training environment. No ToA map is required at deployment. Gradients $\nabla T(\mathbf{x})$ are calculated at grid cell centers via finite differencing and interpolated to provide values in continuous space. The negative gradient then defines the velocity set-point direction $\mathbf{v}_{k}^{\text{set}}$ in the target velocity loss (see Eq. 9), providing a consistent training signal that encourages the robot to make progress toward the goal while avoiding obstacles.
@@ III-D Training
Figure 6: Comparison of attitude control performance without (a) and with (b) the derivative feedback term, $\boldsymbol{\omega}_{d}$, in the attitude controller.
@@ III-E Body Rate Attitude Control
To bridge the sim-to-real gap between point-mass dynamics in simulation and rigid-body dynamics in reality, we adopt a PD attitude controller [13] that tracks desired thrust $\mathbf{F}$, orientation $\mathbf{R}_{d}$, and body rates $\boldsymbol{\omega}_{d}$. The desired rotation $\mathbf{R}_{d}$ aligns the body $\mathbf{z}^{B}$ axis with $\mathbf{F}$ and $\mathbf{x}^{B}$ with the predicted yaw direction $\psi_{d}$. The desired body rates $\boldsymbol{\omega}_{d}$ are computed from ZYX Euler angle rates estimated from consecutive desired attitudes via the Euler angle Jacobian:
@@
Figure 8: Trajectories overlaid on ground-truth point clouds, with a cross-section shown for clarity. Trajectories are colored by speed (red to yellow, up to $v_{\text{max}}=3\ \mathrm{m/s}$) with body frames every 2 s. Start and goal are marked in blue and magenta. Successful trials are marked with a  in Industry 2 (top) and Cave 2 (bottom) environments. ToA maps serve as an inductive bias during training only and are not available during simulation or hardware evaluation.
@@ IV-B Diverse Simulated Environments
We evaluate our planner in both simulated indoor and outdoor environments featuring narrow corridors and large obstacles. The evaluation uses the 11 out-of-distribution scenarios from [14], with depth images rendered by Flightmare [15]. Notably, the policy trained with point-mass dynamics is evaluated under full rigid-body quadrotor dynamics without access to ToA maps. 1,350 simulation trials are conducted across all environments and different start-end location pairs. Further details on the experimental design are available in [14].
We compare our method against two baselines: Back to Newton’s Laws (BNL) [2], trained using the authors’ public codebase [^4] https://github.com/HenryHuYu/DiffPhysDrone , and an ablated version of our model trained without the ToA privileged information (yaw w/o ToA). We do not ablate ToA without yaw prediction, as the ToA gradient may direct the robot toward obstacles outside the depth sensor’s field of view, resulting in unsafe behavior. All approaches are trained using only primitive obstacles and point-mass dynamics. Performance is measured by success rate and failure modes (collisions and timeouts), with success defined as reaching the target without collision within a fixed time limit.
Our method achieves a 86% success rate, 36% higher than the baselines, and has the lowest collision rate across all approaches (Fig. 8). As shown in Fig. 8, The BNL policy struggles to avoid large, flat obstacles due to its fixed heading towards the target. The yaw w/o ToA policy achieves a higher success rate than BNL in environments requiring significant reorientation (e.g., Industry 2), confirming that yaw alignment alone enables the policy to turn around large obstacles; however, without ToA guidance it frequently times out in concave regions. Our full model yaws around large obstacles and learns global navigation cues from the ToA loss, yielding the highest success rate across all environments.
@@ V HARDWARE EXPERIMENTS
Figure 9: Hardware ablation comparing policies trained without LABEL:sub@fig:no_gravity_compensation and with LABEL:sub@fig:gravity_compensation gravity randomization. Top row: flight snapshots with goal setpoint in red. In LABEL:sub@fig:no_gravity_compensation, the start and goal positions coincide, but upon entering autonomous mode, the vehicle quickly loses altitude. With randomization LABEL:sub@fig:gravity_compensation, the robot gains altitude and reaches the goal setpoint. Middle row: plot of z-height from motion capture showing significantly reduced altitude error with the gravity randomized policy. Bottom row: plot of normalized thrust predicted by the policy, where the gravity randomized policy initially outputs $1.3g$, substantially higher than the policy without gravity randomization, compensating for modeling inaccuracies.
A custom quadrotor platform was developed to support hardware experimentation. The vehicle spans $15\text{\,}\mathrm{cm}\text{/}$ from rotor to rotor and is equipped with a forward-facing Intel RealSense D456 depth camera, a downward-facing Lightware SF20/C range finder, and a Matrix Vision BlueFox2 global shutter greyscale camera. All onboard computation is performed by an NVIDIA Orin NX module with 16 GB of RAM. The system has a total mass of $1.7\text{\,}\mathrm{kg}\text{/}$. A TBS Lucid H7 flight controller runs custom Betaflight firmware [^5] https://github.com/betaflight/betaflight to provide IMU data at $1000\text{\,}\mathrm{Hz}\text{/}$. For GPS-denied state estimation, we employ a monocular visual-inertial navigation system developed by Yao [16]. The RealSense camera generates depth images at $60\text{\,}\mathrm{Hz}\text{/}$ with a resolution of 640 $\times$ 480 pixels and the policy runs at $50\text{\,}\mathrm{Hz}\text{/}$.
@@ V-B Outdoor Navigation Experiments
Figure 10: Outdoor obstacle avoidance test under tree canopy (Table III Forest Flight 4). The policy predicts up to $30\text{\,}\mathrm{\SIUnitSymbolDegree}\text{/}$ in yaw to navigate through dense underbrush with speeds up to $3.8\text{\,}\mathrm{m}\text{/}\mathrm{s}$. From top to bottom: VINS trajectory overlaid on terrain map, onboard RGB images, policy depth input (inverted and max pooled), and velocity profile.
@@
Outdoor Flight Arena 1 14.4 12.1 2.0 1.5 0.7
@@ VI CONCLUSION
We developed a simulation and training framework for efficiently learning vision-based navigation policies that map depth and state observations to thrust and heading commands. Trained using simplified point-mass dynamics and privileged information, the policy generalizes to full quadrotor dynamics and reliably navigates through cluttered environments. Domain randomization, especially over gravity, enables robust feedback control, correcting for modeling errors such as a 15% thrust mismatch. Our approach achieves an 86% success rate in simulation and completes 20 hardware flights covering $589\text{\,}\mathrm{m}\text{/}$ without collisions. While the policy generalizes from primitive training obstacles to photorealistic simulation and real-world flights without fine-tuning, it has difficulty with backtracking in maze-like environments (e.g., the Mine scenario) and exhibits initial yaw oscillations. A promising direction for future work is to explore more expressive learned-memory architectures to enhance spatial reasoning and long-horizon planning without relying on explicit maps. Additionally, incorporating task-specific inputs and new objective functions could extend the method’s use to a wider range of applications.
ACKNOWLEDGMENTS